OpenStudy (anonymous):

Two numbers have a sum of 9 and a product of 20. What single-variable quadratic equation in the form ax^2 + bx + c = 0 can be used to represent the product of the two numbers?

OpenStudy (campbell_st):

a + b = 9 (1) ab = 20 (2) looking at equation (1) it can be rewritten as a = 9 - b substitute it into equation (2) b(9 -b) = 9b - b^2 so your quadratic is -b^2 + 9b

OpenStudy (anonymous):

I have a really hard time with questions like these, do you have any tips?

OpenStudy (campbell_st):

nope... just read it carefully the write you equations and try to fins an easy substitution in the question above I forgot to finish it should be -b^2 + 9b - 20 = 0